OpenStudy (anonymous):

Dolls cost $140 per carton and trucks cost $430 per carton. If an order comes in for a total of 100 cartons for $28,500, what was the number of cartons of dolls? (Hint, Let T = Cartons of Trucks)

OpenStudy (anonymous):

140(d)+430(t)=28500 d+t=100 t=100-d 140d+430(100-d)=28500 140d+43000-430d=28500 -290d=-14500 29d=-1450 d=50